A new electric saw for cutting small pieces of lumber in a furniture manufacturing plant has a cost basis of $4000 and a 10-year depreciable life. The estimated SV of the saw is zero at the end of 10 years. Use the DB method to calculate the annual depreciation amount when :

a) R=2/N ( 200% DB Method)
b) R=1.5/N (150% DB Method)

Please use hand calculations with all details.
Answer #1

Computation of Depreciation expense using Declining balance methods:

(a)

When 200% DB method:

Cost of Asset = $4000

Rate of Depreciation =( 2 / N)*100 = (2/10)*100 = 20%

Depreciation expense = Cost of Asset * Depreciation Rate = $4000*20% = $800

(b)

When 150% DB method:

Rate of Depreciation = (1.5/10)*100 = 15%

Depreciation expense = $4000*15% = $600